# FinGAIA: An End-to-End Benchmark for Evaluating AI Agents in Finance

### v1

@@ Abstract
Figure 1: FinGAIA offers an AI Agent evaluation framework for full-process financial operations. From the perspective of business depth, it designs three major scenarios: Basic Business Analysis, Asset Decision Support, and Strategic Risk Control Management. Each primary scenario corresponds to 2, 3, and 3 sub-scenarios respectively, accurately reflecting the diverse business scenarios within the financial industry. The upper part of the image displays the 7 common financial sub-domains covered by the questions FinGAIA has constructed. The middle part illustrates the overall structure of FinGAIA, where business depth increases with the change in color. Concurrently, the demands on the model’s understanding and analytical capabilities regarding financial business operations progressively increase. The side panel shows specific examples corresponding to the sub-scenarios, while the lower part details the specific capabilities required for the AI Agent to solve problems.
Benchmarks Question Type MLD SD MT DE MS FD NOA NoQ
ToolBench MC 5 126,486
API-Bank OE 4 274
WebArena MC 1 812
AgentBench MC 27 1,091
GAIA MC+OE 6 466
FinGAIA (ours) MC+OE 10 407
Table 1: Comparison of Agent Benchmarks. Abbreviations in the header are: MC (Multiple-Choice questions), OE (Open-Ended questions), T/F (True/False questions), NoQ (Number of Questions), NOA (Number of Agents), MLD (Multi-level Difficulty), SD (Scenario Depth), FD (Financial Domains), MT (Multi-tool Support), DE (Dynamic Environment), MS (Multi-step Reasoning).
@@ Introduction
The remainder of this paper is structured as follows. Section 2Related Work reviews related work on AI Agents and Agent benchmarks. Section 3FinGAIA details the construction of FinGAIA, including data collection, question design, and quality control. Sections 4Experiment and 5Results present the experimental setup and results under different difficulty levels, followed by comparative and error analyses. Section 6Conclusion concludes the paper and discusses future directions for financial intelligence agents.
@@
Scenario Depth Financial Scenario Questions
Operational Analytics
Customer Data Analytics 47
Transaction Risk Assessment) 42
All 89
Asset Decision
Financial Data Statistics 101
Loan Credit Analysis 43
Fraud Detection Analysis 41
All 185
Strategic Risk
Risk Management Analysis 42
Portfolio Fund Allocation 40
Market Trend Forecasting 51
All 133
FinGAIA All 407
Table 2: Distribution of Financial Scenarios. This table systematically presents the distribution of FinGAIA tasks across three progressively complex financial scenario categories. Specifically, the Basic Business Analysis category contains 89 tasks, Asset Decision Support includes 185 tasks, and Strategic Risk Management covers 133 tasks, resulting in a total of 407 carefully annotated tasks. This structured presentation enables precise evaluation of the benchmark’s ability to simulate the increasing complexity of real-world financial environments through tiered task design.
@@
In recent years, Agents have evolved from early rule‐based expert systems (Weizenbaum 1966; Barker et al. 1989) into end‐to‐end execution frameworks capable of autonomous learning and multi‐tool invocation. In the late 1990s, data‐driven decision‐making was introduced via collaborative filtering (Resnick et al. 1994) and Q‐Learning (Sutton, Barto et al. 1998). With the advent of deep learning, visual pre‐training (Krizhevsky, Sutskever, and Hinton 2012) and large‐scale language pre‐training (Devlin et al. 2019; Radford et al. 2019) were integrated into Agent architectures, greatly enhancing their multimodal perception and understanding. Methodologically, the “reasoning–execution–feedback” loop proposed by ReAct (Yao et al. 2023) and Toolformer (Schick et al. 2023), together with the closed‐loop toolchain orchestration of Auto-GPT (Significant-Gravitas 2023), endowed Agents with high‐level autonomy in scheduling and invoking multiple tools. At the same time, modular frameworks such as LangChain (Chase 2023) and AgentVerse (Chen et al. 2023) have substantially lowered the barrier to building and deploying multi‐tool Agents.
In the financial domain, Agents have followed a similarly rapid trajectory: beginning with the first Robo-Advisor platforms in 2008 that automated portfolio allocation (Betterment 2008), progressing through the emergence of domain‐specific pre-trained models tailored to financial corpora (Yang, Uy, and Huang 2020), and culminating in recent end-to-end decision systems that integrate real-time market data retrieval, model‐based reasoning, and explainable report generation—thereby driving full-process intelligent financial decision-making (Huang et al. 2024).
Agent Benchmark Analysis
@@
Overall, FinGAIA provides a structured, diverse, and contextually grounded benchmark for evaluating intelligent agents in the financial vertical, enabling comprehensive and objective assessment of their professional competencies, mainly including web browsing capability, document processing capability, multimodal understanding, coding ability, and computational ability. The distribution of tasks across business scenarios and dataset statistics is shown in Table 2, and representative task examples are included in the Appendix 1A.
@@
In the expert review stage, each task was thoroughly inspected by four industry professionals, each with over a decade of financial experience. Review criteria included alignment with real industry practice, logical soundness of the task structure, regulatory compliance, precision of financial terminology, and contextual relevance to the business scenario. A task was accepted into the dataset only if unanimously approved by all four experts, ensuring that each question was well-designed, unambiguous, and reflective of authentic financial operations. Through this rigorous data generation and quality control pipeline, we constructed FinGAIA, a high-quality benchmark for evaluating financial AI agents. The final dataset includes 407 expert-validated tasks. The distribution of tasks across the eight financial scenarios is presented in Table 2.
FinGAIA Question Architecture
The FinGAIA benchmark is built upon representative financial business scenarios and establishes a three-tiered evaluation framework designed to systematically assess the overall performance of intelligent agents in realistic financial environments. The benchmark consists of 407 high-quality tasks, each rigorously validated by experts. Tasks are categorized into three ascending difficulty levels based on task complexity and the depth of required capabilities: Level 1 (Basic Business Analysis), Level 2 (Asset Decision Support), and Level 3 (Strategic Risk Management). These three levels respectively cover 2, 3, and 3 distinct financial scenarios. As the difficulty level increases, both the types of questions and the complexity of tasks intensify, enabling a comprehensive evaluation of an agent’s abilities in financial language understanding, information integration, tool utilization, and complex decision reasoning. Representative examples of each level are provided in the Appendix 1A. Estimates of task difficulty and the average number of tool invocations per scenario are provided in the Appendix Table 5.
Level 1: Basic Business Analysis This level primarily evaluates an agent’s grasp of fundamental financial knowledge and its ability to quickly process multimodal information. Tasks at this level are structurally simple, typically requiring no more than five steps and the use of only one or two tools. The focus is on reading, interpreting, and performing basic calculations over essential financial inputs such as customer data and transaction charts. Suitable for entry-level tasks like customer service or preliminary transaction screening, Level 1 includes two representative scenarios: Customer Data Analytics and Transaction Risk Assessment, comprising a total of 89 tasks.
Level 3: Strategic Risk Management Designed for high-stakes decision-making and advanced risk management, Level 3 features tasks that closely reflect real-world financial practices. These tasks involve a greater number of steps around 10 and require coordinated use of multiple tools, including sequential tool invocation and parameter tuning. Agents are expected to exhibit precise semantic comprehension and integration of financial expertise. This level comprehensively evaluates the agent’s peak performance under complex financial conditions, with a particular focus on accuracy, robustness, and consistency in chain-of-thought reasoning. It includes three core business scenarios: Portfolio Fund Allocation, Risk Management Analysis, and Market Trend Forecasting, comprising 133 tasks.
@@
Table 3: Main Results. The higher the value in the table, the higher the accuracy of the financial AI model. The evaluation assesses performance across three capability tiers: Operational Analytics covering Customer Data Analytics(CDA) and Transaction Risk Assessment(TRA); Asset Decision comprising Financial Data Statistics(FDS), Loan Credit Analysis(LCA), and Fraud Detection Analysis(FDA); and Strategic Risk evaluating Risk Management Analysis(RMA), Portfolio Fund Allocation(PFA), and Market Trend Forecasting(MTF), concluding with the calculation of the Weighted Average (WA) score for each model.
Experiment
Agents
We evaluated 10 currently mainstream Agent frameworks, consisting of 9 closed-source agents and 1 locally deployed open-source framework OWL (CAMEL-AI 2025). 9 closed-source agents were tested via their web interfaces, including ChatGPT DeepResearch (OpenAI 2025), Perplexity DeepResearch (Perplexity AI 2025), Kimi (Moonshot AI 2023), Cashcat DeepResearch (FinStep 2025), Grok DeeperSearch (xAI 2025), Gemini-2.5-pro (Google 2025), Reportify (Beijing Jisha Chengta Technology 2024), and Kompas AI (Kompas AI 2024). And AutoGLM (Zhipu AI 2024) was integrated via a browser plugin. Detailed information is provided in Appendix 2B.
Evaluation Methods
In the actual evaluation process, we designed a differentiated testing process for each Agent’s tool usage capability, aiming to comprehensively assess its performance in end-to-end financial scenarios. For closed-source agents provided as web or application-based services, we conducted interactive tests through manual operation, strictly adhering to the ”zero-shot prompt” paradigm, providing only the problem text and necessary attachments (e.g., Excel files, PDF reports), without offering any examples or guidance. For frameworks that support API calls, we submitted all test questions in bulk through a unified API interface and embedded a preset System Prompt within the request to guide the Agent to output standardized structured answers.
To ensure the accuracy and consistency of the output, we primarily relied on manual review, supplemented by LLM-as-Judge’s automatic interpretation, to perform item-by-item validation of the system’s output. Only when the output results are completely consistent with the standard answer in terms of value, format, and semantics is it considered correct (1 point). Otherwise, it is considered incorrect (0 points). In cases of file processing failure or unanswered questions, they are marked as unassessable (-1 point) to ensure the uniqueness and rigor of the evaluation results.
After completing all data collection and scoring, we computed each agent’s accuracy—both across the full set of tasks and within the Level 1–3 stratified tasks—by dividing its total number of correct responses by the total number of questions. This method ensures a fair and reproducible comparison of each framework’s end-to-end performance and tool-coordination capabilities in financial scenarios.
Results
Main Results
Figure 2: Error types’ distribution across different Agents in FinGAIA tasks.
The performance of financial AI agents varies significantly across different scenarios, demonstrating clear stratification. ChatGPT (DeepResearch) ranked first with an average weighted score (WA) of 48.9%, exhibiting exceptional cross-level stability, particularly excelling in Risk Management Analysis (RMA 60.0) within strategic planning scenarios and Loan Credit Analysis (LCA 58.1%) in tactical advisory scenarios. The second tier, consisting of Perplexity DeepResearch (WA 37%) and Cashcat DeepResearch(WA 33.2%), achieved an average score of 52.4% in Transaction Risk Assessment (TRA) for basic operational tasks, significantly exceeding the industry average by 17.2 points.
Key scenario analysis reveals that market forecasting and fraud detection remain common industry weaknesses, with only ChatGPT demonstrating breakthrough performance in fraud detection (FRA 39.0%). Further analysis indicates that the poorest-performing model, Grok DeeperSearch (FRA 2.4%), suffers from severe cross-modal alignment deficiencies and anomaly pattern recognition blind spots.
Comparative Analysis
As shown in Table 4, unlike Table 3, we calculated the average results across three major scenarios and the overall score to compare human and model performance. The results demonstrate that the best-performing Agent models have surpassed non-experts in all three scenarios and in the composite average score. However, there remains a performance gap of over 32% between current agents and financial experts, indicating that agents still require continuous iterative optimization.
@@ Results
As evidenced by the error distribution analysis, most agents demonstrate relatively balanced error patterns. However, certain models (notably Kimi, ChatGPT(Deep Research), and OWL) exhibit statistically significant error concentrations in specific dimensions: Financial Terminological Bias , Data Type Handling Error , and Entity-Causation Misidentification, respectively. These systematic biases result in disproportionately higher error rates in these categories compared to baseline expectations.Detailed methodological explanations, including sampling procedures and significance testing protocols, along with representative error case studies, are documented in Appendix 2B.
@@ Conclusion
Zero-shot evaluation results show that the best-performing model, ChatGPT (DeepResearch), achieved an overall accuracy of 48.9%. While this performance surpasses that of human finance undergraduates, there remains a significant gap when compared to finance PhDs. This highlights a substantial room for improvement, especially in highly complex, multi-step tasks. Furthermore, our comprehensive error analysis reveals five major capability deficiencies in agents when applied to real-world financial scenarios. As the first agent benchmark rooted in actual financial industry workflows, FinGAIA provides a structured, multi-level evaluation framework for measuring agents’ end-to-end capabilities in financial practice.
@@
Appendix A Details of FinGAIA
@@
We have detailed the FinGAIA data in Table 2. We adopted a three-tiered structure for financial business scenarios, including basic business analysis, asset decision support, and strategic risk management. The detailed information for specific financial business scenarios is presented below.
@@
Appendix B Details of Agents
We list details of the AI Agent evaluated using FinGAIA in Table 7.
@@ Appendix B Details of Agents
Scenario Depth Financial Scenario Tool Count
@@
Customer Data Analytics 2.06
Transaction Risk Assessment) 2.14
@@
Financial Data Statistics 2.63
Loan Credit Analysis 1.98
Fraud Detection Analysis 2.29
@@
Risk Management Analysis 2.02
Portfolio Fund Allocation 2.2
Market Trend Forecasting 2.63
Table 5: The table presents the average number of tool invocations for each financial scenario under the three difficulty levels defined in the FinGAIA dataset. These statistics provide an intuitive reflection of FinGAIA’s progressive task design logic—from simple to complex—facilitating more accurate evaluation of an agent’s performance and adaptability across different financial contexts.
@@
Table 6: The evaluation results of financial AI agents in assessable scenarios.The higher the value in the table, the higher the accuracy of the financial AI agent. The evaluation assesses performance across three capability tiers: Operational Analytics covering Customer Data Analytics(CDA) and Trans-action Risk Assessment(TRA); Asset Decision comprising Financial Data Statistics(FDS), Loan Credit Analysis(LCA), and Fraud Detection Analysis(FDA); and Strategic Risk evaluating Risk Management Analysis(RMA), Portfolio Fund Allocation(PFA), and Market Trend Forecasting(MTF), concluding with the calculation of the Weighted Average (WA) score for each agent. Cashcat DeepResearch is removed since it does not throw error message even for unsupported files to avoid unfair comparison.
@@
Figure 6: A Loan Credit Analysis example where the Agent consults the listed company’s 2024 Annual Report to verify four claims—total shareholders’ equity and its year-on-year change; Q1 vs. Q4 net profit; the identity and stake change of the eighth-largest shareholder; and the signing auditors’ names—and returns any incorrect item numbers, testing data extraction, comparative analysis, and text-matching skills at the Asset Decision Tier.
Figure 7: A Fraud Detection Analysis example where, based on the day’s morning news replay audio, the Agent must 1) identify all companies mentioned under “company developments”; 2) filter for listed companies only; 3) retrieve and compare each company’s stock’s intraday gain at the broadcast time; and 4) output the six-digit code of the top-gaining stock. This tests the Agent’s abilities in audio parsing, entity recognition, data filtering, market-data retrieval, and result generation at the Asset Decision Tier.
Figure 8: This is a Risk Management Analysis (RMA) scenario example, focusing on using Python and the akshare library to fetch daily historical data for the CSI 300, the Hang Seng Index, and the S&P 500 from March 1, 2014 to March 1, 2024, then using the cummax() function to compute each trading day’s historical high and drawdown rate, and finally outputting the maximum drawdown values for the three indices . This tests the Agent’s comprehensive capabilities in long-term market risk assessment and decision support at the Strategic Risk Tier.
Figure 9: This is a Portfolio Fund Allocation scenario example, focusing on dynamic parameter mapping and portfolio holding analysis—the Agent must first identify the number of animal species X shown in the image, then rank funds by their drawdowns from April 1, 2021 to April 1, 2022, select the fund with the Xth largest drawdown, retrieve that fund’s 2021 holdings, identify the security with the highest net‐asset proportion, and output only the numeric part of its security code. This task tests the Agent’s integrated abilities in visual information extraction, dynamic metric mapping, fund performance ranking, and portfolio holding analysis, illustrating the Strategic Risk Tier’s application value for asset allocation optimization and return forecasting.
Figure 10: This is a Market Trend Forecasting (MKT: Market Trend Forecasting) scenario example, focusing on a multi‐security screening task based on up/down pattern matching—the Agent must identify the target security’s five‐day rise/fall sequence (hollow red bars indicate up days; solid cyan bars indicate down days), then compare that sequence against securities 000001- 000007 over the five trading days from February 5 to February 11, 2025, selecting those whose patterns match exactly and returning their six‐digit codes. This task tests the Agent’s integrated abilities in time‐series pattern recognition, pattern matching, and batch screening of multiple securities, illustrating the Strategic Risk Tier’s value in historical trend benchmarking and forecasting.

### v2

# FinGAIA: A Chinese Benchmark for AI Agents in Real-World Financial Domain

@@ Abstract
Figure 1: FinGAIA offers an AI Agent evaluation framework for full-process financial operations. From the perspective of business depth, it features three major scenarios: Operational Analytics, Asset Decision, and Strategic Risk. Each primary scenario includes 2, 3, and 3 sub-scenarios, respectively, accurately reflecting the diverse business needs of the financial industry. The upper part of the image shows the 7 common financial sub-domains covered by FinGAIA’s constructed questions. The left section outlines FinGAIA’s overall structure, where business depth increases as the color gradient darkens. Simultaneously, the framework progressively raises demands on the model’s understanding and analytical capabilities regarding financial operations.The lower left part details the specific domains and competencies required for the AI Agent to solve problems, while the right panel provides a concrete example, including the question, steps, and answer.
Benchmarks Question Type Multi-level Difficulty Scenario Depth Multi-tool Support Dynamic Environment Financial Domains
ToolBench MC
API-Bank OE
WebArena MC
AgentBench MC
GAIA MC+OE
FinGAIA (ours) MC+OE
Table 1: Comparison of Agent Benchmarks. Question Type abbreviations: MC (Multiple-Choice questions), OE (Open-Ended questions).
@@ Introduction
In recent years, Agent technology has evolved from basic multi-step tool invocation to autonomous end-to-end systems. Initially, frameworks like LangChain proposed modular multi-step reasoning and external tool chaining mechanisms (Chase 2023). Subsequently, Toolformer enabled Agents to autonomously determine and invoke appropriate tools through self-supervision, significantly enhancing decision-making efficiency (Schick et al. 2023). ReAct integrated reasoning with action decision-making, enabling real-time API calls and feedback (Yao et al. 2023). Building on this, open-source projects like AutoGPT and BabyAGI introduced a ”decompose-execute-feedback” loop, achieving end-to-end adaptive optimization (Significant-Gravitas 2023; Nakajima 2023). Concurrently, OpenAI’s Function Calling and Plugins ecosystem standardized interfaces for secure and efficient interaction between Agents and third-party services (OpenAI 2023b, a). Academia has also provided solutions in multi-model collaboration, task planning, and operational monitoring, offering systematic support for Agent development and operation (Shen et al. 2023; Prasad et al. 2023; Smith 2024). In the financial domain, Agents have undergone rapid iteration: from their initial form as digital advisors (Robo-advisors), providing algorithmic asset allocation for individual investors with low-cost, automated portfolio management services (Fisch, Laboure, and Turner 2019); to domain-specific pre-training systems for financial corpora (Yang, Uy, and Huang 2020); and further to integrated real-time market data capture, model inference, and interpretable report generation, driving the intelligent automation of financial decision-making processes (Huang et al. 2024).
Agent Benchmark History
@@
Overall, FinGAIA provides a structured, diverse, and contextually grounded benchmark for evaluating intelligent agents in the financial vertical, enabling comprehensive and objective assessment of their professional competencies, mainly including web browsing capability, document processing capability, multimodal understanding, coding ability, and computational ability. The distribution of tasks across business scenarios and dataset statistics is shown in Appendix Table 4, and representative task examples are included in the Appendix A.
@@
In the expert review stage, each task was thoroughly inspected by four industry professionals, each with over a decade of financial experience. Review criteria included alignment with real industry practice, logical soundness of the task structure, regulatory compliance, precision of financial terminology, and contextual relevance to the business scenario. A task was accepted into the dataset only if unanimously approved by all four experts, ensuring that each question was well-designed, unambiguous, and reflective of authentic financial operations. Through this rigorous data generation and quality control pipeline, we constructed FinGAIA, a high-quality benchmark for evaluating financial AI agents. The final dataset includes 407 expert-validated tasks.
@@
Table 2: Main Results. The higher the value in the table, the higher the accuracy of the financial AI model. The evaluation assesses performance across three capability tiers: Operational Analytics covering Customer Data Analytics(CDA) and Transaction Risk Assessment(TRA); Asset Decision comprising Financial Data Statistics(FDS), Loan Credit Analysis(LCA), and Fraud Detection Analysis(FDA); and Strategic Risk evaluating Risk Management Analysis(RMA), Portfolio Fund Allocation(PFA), and Market Trend Forecasting(MTF), concluding with the calculation of the Weighted Average (WA) score for each model. The highest score in each scenario is bolded in the table.
FinGAIA Question Architecture
The FinGAIA benchmark is built upon representative financial business scenarios and establishes a three-tiered evaluation framework designed to systematically assess the overall performance of intelligent agents in realistic financial environments. The benchmark consists of 407 high-quality tasks, each rigorously validated by experts. Tasks are categorized into three ascending difficulty levels based on task complexity and the depth of required capabilities: Level 1 (Basic Business Analysis), Level 2 (Asset Decision Support), and Level 3 (Strategic Risk Management). These three levels respectively cover 2, 3, and 3 distinct financial scenarios. As the difficulty level increases, both the types of questions and the complexity of tasks intensify, enabling a comprehensive evaluation of an agent’s abilities in financial language understanding, information integration, tool utilization, and complex decision reasoning. Representative examples of each level are provided in the Appendix A. The average number of tool invocations per scenario are provided in Table 4.
Level 1: Basic Business Analysis This level primarily evaluates an agent’s grasp of fundamental financial knowledge and its ability to quickly process multi-modal information. Tasks at this level are structurally simple, typically requiring no more than five steps and the use of only one or two tools. The focus is on reading, interpreting, and performing basic calculations over essential financial inputs such as customer data and transaction charts. Suitable for entry-level tasks like customer service or preliminary transaction screening, Level 1 includes two representative scenarios: Customer Data Analytics and Transaction Risk Assessment, comprising a total of 89 tasks.
Level 3: Strategic Risk Management Designed for high-stakes decision-making and advanced risk management, Level 3 features tasks that closely reflect real-world financial practices. These tasks involve a greater number of steps around 10 and require coordinated use of multiple tools, including sequential tool invocation and parameter tuning. Agents are expected to exhibit precise semantic comprehension and integration of financial expertise. This level comprehensively evaluates the agent’s peak performance under complex financial conditions. It includes three core business scenarios: Portfolio Fund Allocation, Risk Management Analysis, and Market Trend Forecasting, comprising 133 tasks.
@@
Experiment
Agents
We evaluated 10 currently mainstream Agent frameworks, consisting of 9 closed-source agents and 1 locally deployed open-source framework OWL (CAMEL-AI 2025). 9 closed-source agents were tested via their web interfaces, including ChatGPT DeepResearch (OpenAI 2025), Perplexity DeepResearch (Perplexity AI 2025), Kimi (Moonshot AI 2023), Cashcat DeepResearch (FinStep 2025), Grok DeeperSearch (xAI 2025), Gemini-2.5-pro (Google 2025), Reportify (Beijing Jisha Chengta Technology 2024), and Kompas AI (Kompas AI 2024). And AutoGLM (Zhipu AI 2024) was integrated via a browser plugin. Detailed information is provided in Appendix B.
Figure 2: This radar chart compares error distributions across six financial AI agents in FinGAIA evaluations, with vertex position reflecting severity in five error types. Axis scales (0–0.5) quantify performance deviations for targeted improvement.
Evaluation Methods
In the actual evaluation process, we employed differentiated testing protocols tailored to each agent’s tool‑invocation capabilities. For closed‑source agents accessed via web or application interfaces, we conducted manual zero‑shot interaction tests—providing only the task prompt and necessary attachments (e.g., Excel files, PDF reports) without any examples or guidance. For API‑enabled frameworks, we submitted all test items in batch through a unified API endpoint, embedding a predefined system prompt to elicit standardized, structured outputs.
During the evaluation, we recorded each agent’s score for correctly executed tasks across all test scenarios. To ensure the accuracy and consistency of the output, we primarily relied on manual review, supplemented by LLM-as-Judge’s automatic interpretation (Xu et al. 2023), to perform item-by-item validation of the system’s output.
After completing all data collection and scoring, we computed each agent’s accuracy—both across the full set of tasks and within the Level 1–3 stratified tasks—by dividing its total number of correct responses by the total number of questions. This method ensures a fair and reproducible comparison of each framework’s end-to-end performance and tool-coordination capabilities in financial scenarios.In Appendix B, we excluded tasks unprocessable due to unsupported file formats and recalculated accuracy on the remaining successful tasks to illustrate each agent’s performance in executable scenarios.
Results
Main Results
The performance of financial AI agents varies significantly across different scenarios, demonstrating clear stratification. ChatGPT (DeepResearch) ranked first with an average weighted score (WA) of 48.9, exhibiting exceptional cross-level stability, particularly excelling in Risk Management Analysis (RMA 60.0) within strategic planning scenarios and Loan Credit Analysis (LCA 58.1) in tactical advisory scenarios. The second tier, consisting of Perplexity DeepResearch (WA 37) and Cashcat DeepResearch(WA 33.2), achieved an average score of 52.4 in Transaction Risk Assessment (TRA) for basic operational tasks, significantly exceeding the industry average by 17.2 points.
Key scenario analysis reveals that market forecasting and fraud detection remain common industry weaknesses, with only ChatGPT demonstrating breakthrough performance in fraud detection (FRA 39.0). Further analysis indicates that the poorest-performing model, Grok DeeperSearch (FRA 2.4), suffers from severe cross-modal alignment deficiencies and anomaly pattern recognition blind spots.
Comparative Analysis
As shown in Table 3, unlike Table 2, we calculated the average results across three major scenarios and the overall score to compare human and model performance. The results demonstrate that the best-performing Agent models have surpassed non-experts in all three scenarios and in the composite average score. However, there remains a performance gap of over 32 between current agents and financial experts, indicating that agents still require continuous iterative optimization.
@@ Results
As evidenced by the error distribution analysis, most agents demonstrate relatively balanced error patterns. However, certain models (notably Kimi, Reportify, and Cashcat DeepResearch) exhibit statistically significant error concentrations in specific dimensions: Operational Process Awareness Barrier, Hallucinatory Financial Reasoning, and Entity-Causation Misidentification, respectively. These systematic biases result in disproportionately higher error rates in these categories compared to baseline expectations.Detailed methodological explanations, including sampling procedures and significance testing protocols, along with representative error case studies, are documented in Appendix B.
@@ Conclusion
Zero-shot evaluation results show that the best-performing model, ChatGPT (DeepResearch), achieved an overall accuracy of 48.9. While this performance surpasses that of human finance undergraduates, there remains a significant gap when compared to finance PhDs. This highlights a substantial room for improvement, especially in highly complex, multi-step tasks. Furthermore, our comprehensive error analysis reveals five major capability deficiencies in agents when applied to real-world financial scenarios. As the first agent benchmark rooted in actual financial industry workflows, FinGAIA provides a structured, multi-level evaluation framework for measuring agents’ end-to-end capabilities in financial practice.
@@
Appendix A A Details of FinGAIA
To promote transparency and facilitate the reproducibility of our findings, a portion of the FinGAIA benchmark data is provided. The data is hosted anonymously at: https://github.com/SUFE-AIFLM-Lab/FinGAIA
@@ Appendix A A Details of FinGAIA
We detail the FinGAIA data in Table 4. We adopted a three-tiered structure for financial business scenarios, including basic business analysis, asset decision support, and strategic risk management. The detailed information for specific financial business scenarios is presented below.
@@
Appendix B B Details of Agents
We list details of the AI Agent evaluated using FinGAIA in Table 6, where tasks unprocessable due to unsupported file formats have been excluded.
@@ Appendix B B Details of Agents
Scenario Depth Financial Scenario Questions Tool Count
@@
Customer Data Analytics 47 2.06
Transaction Risk Assessment 42 2.14
All 89
@@
Financial Data Statistics 101 2.63
Loan Credit Analysis 43 1.98
Fraud Detection Analysis 41 2.29
All 185
@@
Risk Management Analysis 42 2.02
Portfolio Fund Allocation 40 2.2
Market Trend Forecasting 51 2.63
All 133
FinGAIA All 407
Table 4: Distribution of Financial Scenarios. This table systematically presents the distribution of tasks across three progressively complex financial scenario categories in the FinGAIA dataset—Basic Business Analysis, Asset Decision Support, and Strategic Risk Management—along with the average number of tool invocations for each scenario. The average tool invocation statistics further provide an intuitive reflection of FinGAIA’s tiered task design logic, facilitating more accurate evaluation of an agent’s task execution capabilities and tool adaptability across financial scenarios of varying complexity.
@@
Table 5: The evaluation results of financial AI agents in assessable scenarios.The higher the value in the table, the higher the accuracy of the financial AI agent. The evaluation assesses performance across three capability tiers: Operational Analytics covering Customer Data Analytics(CDA) and Trans-action Risk Assessment(TRA); Asset Decision comprising Financial Data Statistics(FDS), Loan Credit Analysis(LCA), and Fraud Detection Analysis(FDA); and Strategic Risk evaluating Risk Management Analysis(RMA), Portfolio Fund Allocation(PFA), and Market Trend Forecasting(MTF), concluding with the calculation of the Weighted Average (WA) score for each agent. Cashcat DeepResearch is removed since it does not throw error message even for unsupported files to avoid unfair comparison.The highest score in each scenario is bolded in the table.
@@
Figure 6: A Loan Credit Analysis scenario example where the Agent consults the listed company’s 2024 Annual Report to verify four claims—total shareholders’ equity; Q1 vs. Q4 net profit; the identity and stake change of the eighth-largest shareholder; and the signing auditors’ names—and returns any incorrect item numbers, testing data extraction, comparative analysis, and text-matching skills at the Asset Decision Tier.
Figure 7: A Fraud Detection Analysis scenario example where, based on the day’s morning news replay audio, the Agent must 1) identify all companies mentioned under “company developments”; 2) filter for listed companies only; 3) retrieve and compare each company’s stock’s intraday gain at the broadcast time; and 4) output the six-digit code of the top-gaining stock. This tests the Agent’s abilities in audio parsing, entity recognition, data filtering, market-data retrieval, and result generation at the Asset Decision Tier.
Figure 8: This is a Risk Management Analysis scenario example, focusing on using Python and the akshare library to fetch daily historical data for the CSI 300, the Hang Seng Index, and the S&P 500 from March 1, 2014 to March 1, 2024, then using the cummax() function to compute each trading day’s historical high and drawdown rate, and finally outputting the maximum drawdown values for the three indices . This tests the Agent’s comprehensive capabilities in long-term market risk assessment and decision support at the Strategic Risk Tier.
Figure 9: This is a Portfolio Fund Allocation scenario example, focusing on dynamic parameter mapping and portfolio holding analysis—the Agent must first identify the number of animal species X shown in the image, then rank funds by their drawdowns, select the fund with the Xth largest drawdown, retrieve that fund’s 2021 holdings, identify the security with the highest net‐asset proportion. This task tests the Agent’s integrated abilities in visual information extraction, dynamic metric mapping, fund performance ranking, and portfolio holding analysis, illustrating the Strategic Risk Tier’s application value for asset allocation optimization and return forecasting.
Figure 10: This is a Market Trend Forecasting scenario example, focusing on a multi‐security screening task based on up/down pattern matching—the Agent must identify the target security’s five‐day rise/fall sequence, then compare that sequence against securities 000001- 000007 over the five trading days, selecting those whose patterns match exactly and returning their six‐digit codes. This task tests the Agent’s integrated abilities in time‐series pattern recognition, pattern matching, and batch screening of multiple securities, illustrating the Strategic Risk Tier’s value in historical trend benchmarking and forecasting.